How much water should we add to 20 kg of seawater to change its concentration from 3% to 2%?

Answer by Boreal(15235)

x=amount of water in kg. (0 salt)
start with 20 kg at .03 (3%) (salt)=.60
add x kg of water with 0 salt.
will have (20 + x) kg of water with 2 % salt.
Distribute. 02* (20+x)=0.4 +0.02x
This will equal .60, the desired concentration.
So, 0.4 +0.02x=.60
Multiply the whole equation by 100. Not necessary, but will be easier to work with
40 + 2 x=60
Subtract 40 from both sides
2x=20
Divide by 2 on both sides.
x=10 liters.
Check. We are decreasing the concentration 1/3.
We add 10 liters of water to 20, and the final amount, 30, has 1/3 less salt, and 1/3 of it was without salt.